a. What is ATC for each level of output listed in the table? Output $TC $ATC...

a.

What is ATC for each level of output listed in the table?

Output $TC $ATC

75,000 $100,000 $       [ Select ]    ["1.33", "1", "1.20", "0.67"]      

100,000 $125,000 $ [ Select ]       ["1.10", "1", "1.25", "0.87"]      

125,000 $140,000 $1.12

150,000 $150,000 $1.00

b. Is this a decreasing-cost industry?    [ Select ] ["no", "need more information", "yes"]      

c. Suppose that the price of a bottle of vitamins is $1.33 and that at that price the total quantity demanded by consumers is 75,000 bottles.

     How many firms will there likely be in this industry?     [ Select ]      ["none", "many", "1"]          firm(s).

d. Given the information above, part  c, what kind of competition would you expect this to be [ Select ]    ["oligopoly", "monopolistic competition", "monopoly"]      
Answer #1

What is ATC for each level of output listed in the table?

Quantity 75000

Total cost :1,00,000

ATC=TC/Q

=1.33

Quantity :100000

Total cost :1,25,000

ATC=TC / Q

=1,25,000/100000

=1.25.

b. Is this a decreasing-cost industry?

Yes.

Explanation:

It is decreasing industry as with output increasing average total cost is decreasing.

c. Suppose that the price of a bottle of vitamins is $1.33 and that at that price the total quantity demanded by consumers is 75,000 bottles.

How many firms will there likely be in this industry?

Answer : many

Explanation : when price is equals to average total cost we can assume that there are many firms as in long-run firm makes zero economic profit.

d. Given the information above, part c, what kind of competition would you expect this to be

monopolistic competition

Explanation :

Monopolistic competitive firm earns zero economic profit in long run. Monopoly and oligopoly firms can earn positive profits. So it is more likely to monopolistic competition.